# Local Differential Privacy for Tensors in Distributed Computing Systems

### v2

@@ III-B Randomized Response
It satisfies $\ln(\frac{p}{1-p})$-LDP.
IV Local Differential Privacy of Tensors
This section presents the definition of LDP for tensor-valued data, followed by details of the proposed TLDP and weighted TLDP, including their motivation, theorem, and proof.
@@ IV Local Differential Privacy of Tensors
$\forall\mathcal{X}^{\ast}\in Range(\textbf{M}),Pr[\textbf{M}(\mathcal{X})=\mathcal{X}^{\ast}]\leq e^{\epsilon}Pr[\textbf{M}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}],$
@@
In Local Differential Privacy (LDP), users are unaware of each other’s records. As a result, the concept of global sensitivity, which quantifies the maximum change in query results between two adjacent datasets in differential privacy, does not apply. In this paper, we use $\Delta$ to denote the range of variation of the raw data. It is important to note that $\Delta$ differs from global sensitivity: the former is only related to the users’ initial data, while global sensitivity pertains to query functions and is affected by the size of tensors.
Comparing Definition 1 with Definition 2, it is evident that Definition 1 is a special case of Definition 2, where the tensor order is one (i.e., $N=1,I_{1}=1$).
IV-A TLDP
@@
Originally designed for scalar data with binary outcomes, the randomized response technique has been extended to multi-valued scenarios by encoding scalars into higher-dimensional vectors or adjusting flipping probabilities. However, applying these methods to tensor data presents unique challenges. Encoding each tensor element individually would require substantial storage resources, exponentially increasing algorithmic complexity. Conversely, given our research’s need to retain information from each tensor rather than relying solely on aggregation, methods that randomly perturb selected tensor elements could significantly reduce the utility of the original data.
In light of these challenges, there is a clear need for more effective methods that balance privacy protection and the utility of tensor data. Building upon the aforementioned concepts, we propose a novel approach named TLDP for achieving tensor differential privacy by selectively perturbing specific positions with noise determined via randomized response. In this method, each tensor element remains its original value with a probability of $p$ and introduces calibrated noise (either Laplace or Gaussian) with a probability of $1-p$. Our method ensures privacy protection for each user and demonstrates that compared to directly adding a noise tensor, the inherent randomness of randomized response allows for meeting privacy budgets without requiring excessive noise, thereby maintaining a balance between privacy preservation and data utility.
@@
    $\leq\frac{p}{1-p}\cdot 2b\cdot e^{\frac{I\Delta}{b}}=\frac{e^{\epsilon-\frac{I\Delta}{b}}}{2b}\cdot 2b\cdot e^{\frac{I\Delta}{b}}=e^{\epsilon}.$
  2. 2.
    Gaussian noise
@@
Although DP and LDP differ in several respects, when both perturbation methods involve injecting Laplace noise or Gaussian noise, the amount of noise under the same $\epsilon$ can reflect the impact of each on the accuracy of the results.
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{2b_{1}})^{I}}{(\frac{1}{2b_{1}}\cdot e^{-\frac{\Delta}{b_{1}}})^{I}}$ (5)
    $=e^{\frac{I\Delta}{b_{1}}}.$
@@
    Combining with the feature of Laplace distribution that the larger the scale parameter $b$, the greater the noise, we conclude that TLDP injects less Laplace noise while guaranteeing the same privacy budget than adding noise tensor directly. In particular, the larger the size of the data, the more significant the noise reduction effect will be.
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{\sigma_{1}\sqrt{2\pi}})^{I}}{(\frac{1}{\sigma_{1}\sqrt{2\pi}}\cdot e^{-\frac{\Delta^{2}}{2\sigma_{1}^{2}}})^{I}}$ (7)
    $=e^{\frac{I\Delta^{2}}{2\sigma_{1}^{2}}}.$
@@
IV-B Weighted TLDP
The overzealous pursuit of privacy protection often results in significant compromises to data usability. Striking a balance between these two facets is crucial, especially considering that data usability is not a variable we can easily manipulate. Therefore, attention must be focused on fine-tuning the privacy protection algorithms. Our innovative proposal involves assigning specific weights to the tensor components, where a higher weight indicates greater sensitivity at that particular position, consequently reducing the likelihood of its preservation. To simplify the algorithm, we introduce a weight matrix into the tensor data to be processed rather than using a weight tensor. For example, a color image is a three-order tensor of size $A\times B\times 3$, and a weight matrix of size $A\times B$ is assigned based on the image content. Upon analyzing the following proof, it is evident that, given the same $\epsilon$, the presence of the weight matrix reduces the probability of retaining the original data compared to Algorithm 1, thereby providing stronger privacy protection.
@@ IV-B Weighted TLDP
    $=\frac{(1-w_{mn})e^{\epsilon-\frac{I\Delta}{b}}}{2b+w_{mn}e^{\epsilon-\frac{I\Delta}{b}}}\cdot 2b\cdot e^{\frac{I\Delta}{b}}$
    $\leq\frac{(1-w_{mn})e^{\epsilon-\frac{I\Delta}{b}}}{2b}\cdot 2b\cdot e^{\frac{I\Delta}{b}}$
    $=(1-w_{mn})e^{\epsilon}<e^{\epsilon}.$
@@
    $N(x|\sigma^{2})=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}},$ (11)
@@
    $=\frac{(1-w_{mn})e^{\epsilon-\frac{I\Delta^{2}}{2\sigma^{2}}}}{\sigma\sqrt{2\pi}+w_{mn}e^{\epsilon-\frac{I\Delta^{2}}{2\sigma^{2}}}}\cdot\sigma\sqrt{2\pi}\cdot e^{\frac{I\Delta^{2}}{2\sigma^{2}}}$
    $\leq\frac{(1-w_{mn})e^{\epsilon-\frac{I\Delta^{2}}{2\sigma^{2}}}}{\sigma\sqrt{2\pi}}\cdot\sigma\sqrt{2\pi}\cdot e^{\frac{I\Delta^{2}}{2\sigma^{2}}}$
    $=(1-w_{mn})e^{\epsilon}<e^{\epsilon}.$
V Evaluations
@@
V-A Setup
V-A1 Datasets and Tasks
@@
For the computer vision task, we select MNIST [36], CIFAR-10 [37], and SVHN [38] as the datasets. MobileNet (pre-trained on ImageNet)  [^1] https://pytorch.org/vision/main/models/generated/torchvision.models.mobilenet_v2.html#torchvision.models.MobileNet_V2_Weights is used for the computer vision task. The model is trained with the Adam optimizer with a learning rate of 0.001. It uses the cross-entropy loss and has a batch size of 64. We train it for 10 epochs. The MNIST dataset consists of 60,000 training samples and 10,000 test samples of 28$\times$28 pixel grayscale images of handwritten digits. CIFAR-10 contains 50,000 training samples and 10,000 test samples and has 10 classes of common objects, e.g., birds, airplanes, and cats. SVHN includes door number images collected from Google Street View, and it has 73,257 training images and 26,032 testing images.
@@
V-A2 Baselines and Metrics
We compare TLDP with other differential privacy mechanisms designed for high-dimensional data. The baselines contain: the Laplace mechanism (Laplace), Gaussian mechanism (Gaussian), Matrix-Variate Gaussian (MVG) [16], and Independent Directional Noise (IDN) [17]. We implement TLDP-L, TLDP-G, TLDP-L-w, and TLDP-G-w as solutions for applying local differential privacy on private data (see Section IV), where TLDP-L and TLDP-G denote TLDP that uses the Laplace and Gaussian noise while TLDP-L-w and TLDP-G-w represent TLDP when using Laplace or Gaussian noise while applying a weight matrix for improving utility. Unlike previous studies where noise designed for scalar-valued queries is directly applied to tensors, we mathematically prove that the Laplace and Gaussian noise we add satisfies LDP for tensors (Definition 2), as detailed in Theorem 2. IDN, on the other hand, is a $(\epsilon,\delta)$-differential privacy mechanism for tensor-valued queries. In all experiments, the utility subspace $W$ is set to the identity matrix $E$.
@@ V-A2 Baselines and Metrics
$F1_{i}=2\times\frac{P_{i}\times R_{i}}{P_{i}+R_{i}},\$ (15)
@@
V-A3 Privacy-Preserving Targets
In our experiments, we simulate the distributed computation system using three users and one central server. The users compute their local models using their private dataset and transmit the trained model parameters to the central server. The central server aggregates the received model parameters, computes an average of them, and sends it back to all the users. We denote the models trained with each user’s dataset as local models and the aggregated model on the central server as the global model. Here, we consider three typical tasks in the distributed computation system, i.e., original data, training features, and model parameters. For the first type, users simply transmit their private data to the server, leveraging the enormous computing resources of the server to process and analyze the data. The second type focuses on features extracted from private datasets. This scenario frequently arises in distributed model training across multiple parties, where users process raw data to extract features and transmit the extracted feature embeddings instead of raw data to the central server for model training to avoid privacy leakage. The query mechanism outputs intermediate data/features with differential privacy guarantees, which are then utilized for subsequent training processes on the server side. The third type emphasizes the protection of model parameters. Sharing model parameters is very common in distributed systems, especially when federated learning [41] is evolved for model computation without the need for direct data or feature sharing between data owners and a parameter server. However, research [42] has found that attackers could easily infer the original data used to train the model from intercepted model parameters. Given that the trained model parameters are sensitive, the query function returns a differentially private version of them before transmitting. We present the implementation details categorized by type as follows.
V-B Implementation Details
V-B1 Type I: Private Testing Data
The datasets used in Type I include MNIST, CIFAR-10, and SVHN. The setup is illustrated in Table II. In this type, the training data is likely to be private and should be preserved. Users manipulate their respective data and send it directly to a central server, which uses this information to train the model parameters before transmitting the updated parameters back. No preprocessing or size-trimming is applied to the raw data, ensuring that the accuracy of the experimental results is maintained at the highest possible level.
Additionally, for the MNIST dataset, we include comparisons with the TLDP-L-w and TLDP-G-w methods. In these approaches, the weights are determined based on the grayscale values of the images: higher grayscale values correspond to higher weights. This setting is reasonable since, as the dataset has been preprocessed, apart from the digits with values greater than 0, all other regions are set to 0.
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2\sqrt{I_{1}I_{2}\dots I_{N}},$ (17)
since the range of data is set to $(-1,1)$.
TABLE II: Datasets and parameter settings for Type I
Dataset MNIST Cifar-10 SVHN
Model MobileNet MobileNet MobileNet
Training Data 60000 50000 73257
Testing Data 10000 10000 26032
Batch size 64 64 64
$\Delta$ 256 256 256
V-B2 Type II: Private Training Features
For Type II, we select MNIST, CIFAR-10, and SVHN as experimental datasets. The setup is shown in Table III. DarkNet and IMDB are not used in this task due to the absence of pre-trained models, and without them, the extracted feature embeddings are prone to be less meaningful. We consider the features extracted from the original training data of these datasets to be confidential and are committed to protecting privacy. To enhance model performance, we substitute the activation function in each layer of MobileNet from ReLU6 to tanh()($\cdot$)( ⋅ ), ensuring that the output features are normalized within the range of $(-1,1)$. MobileNet is divided into two parts: the feature extractor (i.e., all layers except the final fully connected layer) and the classifier (i.e., the final fully connected layer). After extracting features from the feature extractor, noise is added to the features, and only the classifier is trained and tested.

### v3

@@ 3.2 Randomized Response
Therefore, it satisfies $\ln(\frac{p}{1-p})$-LDP.
4 Local Differential Privacy of Tensors
In this section, we will formally introduce the proposed local differential privacy method for tensors. First, we present the unified definition of LDP for tensor-valued data, then the motivation, theorem, and proof of TLDP and weighted TLDP are presented.
@@ 4 Local Differential Privacy of Tensors
$\forall\mathcal{X}^{\ast}\in Range(\textbf{M}),Pr[\textbf{M}(\mathcal{X})=\mathcal{X}^{\ast}]\leq e^{\epsilon}\cdot Pr[\textbf{M}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}],$
@@
[…]
Comparing Definition 1 with Definition 2, it is evident that Definition 1 is a special case of Definition 2. Specifically, when the tensor order is reduced to one, the multi-dimensional tensor structure in Definition 2 degenerates into a one-dimensional vector form, which aligns exactly with the data representation specified in Definition 1. In contrast, Definition 2 relaxes the constraint on tensor order, allowing $N$ to take any positive integer value and supporting multi-indexed tensor inputs (e.g., $N=2$ for matrix-valued data and $N=3$ for 3D tensor data) by introducing the index set $\{I_{1},I_{2},...,I_{N}\}$. This hierarchical relationship between the two definitions is not only theoretically consistent but also practically valuable: it means that any conclusion derived under Definition 1 can be naturally extended to the broader framework of Definition 2, while Definition 2 retains the ability to handle simpler one-dimensional scenarios through its special case.
4.1 TLDP
@@
Originally designed for scalar data with binary outcomes, the randomized response technique has been extended to multi-valued scenarios by encoding scalars into higher-dimensional vectors or adjusting flipping probabilities. However, applying these methods to tensor data presents unique challenges. Encoding each tensor element individually would require substantial storage resources, exponentially increasing algorithmic complexity. Conversely, given our research’s need to retain information from each tensor rather than relying solely on aggregation, methods that randomly perturb selected tensor elements could significantly reduce the utility of the original data. Both refinement strategies are unable to meet our requirements of perturbing the raw tensors as little as possible while preserving differential privacy and avoiding other limiting factors such as excessive storage space.
In light of these challenges, there is a clear need for more effective methods that balance privacy protection and the utility of tensor data. Building upon the aforementioned concepts, we propose a novel approach named TLDP for achieving tensor differential privacy by injecting noise into specific positions of tensors, which are determined via randomized response. In this method, each tensor element retains its original value with a probability of $p$ and introduces calibrated noise (either Laplace or Gaussian noise) with a probability of 1p1-p. The randomized response mechanism still determines whether a piece of data will be altered, except that the way of alteration is no longer flipping but injecting noise. The overall procedure is shown in Algorithm 1. Our method ensures privacy protection for each user and demonstrates that, compared to directly adding a noise tensor, the inherent randomness of randomized response allows for meeting privacy budgets without requiring excessive noise, thereby maintaining a balance between privacy preservation and data utility. In the following, we will employ rigorous mathematical proofs to verify the privacy preservation property of this algorithm.
@@
$\leq\frac{p}{1-p}\cdot 2b\cdot e^{\frac{I\Delta}{b}}=\frac{e^{\epsilon-\frac{I\Delta}{b}}}{2b}\cdot 2b\cdot e^{\frac{I\Delta}{b}}$
$=e^{\epsilon}.$
(2) Gaussian noise
@@
The penultimate equations in the above proofs are derived by substituting the formula for $p$ from Algorithm 1. The same logic applies to subsequent proofs, and no further explanation will be provided hereafter.
Although DP and LDP differ in several respects, when both perturbation methods involve injecting Laplace noise or Gaussian noise, the amount of noise under the same $\epsilon$ can reflect the impact of each on the accuracy of the results. The following Theorem 2 demonstrates that, under the same privacy budget, the TLDP method can significantly reduce the amount of noise injected compared to adding noise to all components.
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{2b_{1}})^{I}}{(\frac{1}{2b_{1}}\cdot e^{-\frac{\Delta}{b_{1}}})^{I}}=e^{\frac{I\Delta}{b_{1}}}.$ (5)
@@
    Combining with the feature of Laplace distribution that the larger the scale parameter $b$, the greater the noise, we conclude that TLDP injects less Laplace noise while guaranteeing the same privacy budget as adding noise tensor directly. In particular, the larger the size of the data, the more significant the noise reduction effect will be.
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{\sigma_{1}\sqrt{2\pi}})^{I}}{(\frac{1}{\sigma_{1}\sqrt{2\pi}}\cdot e^{-\frac{\Delta^{2}}{2\sigma_{1}^{2}}})^{I}}=e^{\frac{I\Delta^{2}}{2\sigma_{1}^{2}}}.$ (7)
@@
4.2 Weighted TLDP
The overzealous pursuit of privacy protection often results in significant compromises to data usability. Striking a balance between these two facets is crucial, especially considering that data usability is not a variable we can easily manipulate. Therefore, attention must be focused on fine-tuning the privacy protection algorithms. Our innovative proposal involves assigning specific weights to the tensor components, where a higher weight indicates greater sensitivity at that particular position, consequently reducing the likelihood of its preservation. To simplify the algorithm, we introduce a weight matrix into the tensor data to be processed rather than using a weight tensor. For example, a color image is a three-order tensor of size $A\times B\times 3$, and a weight matrix of size $A\times B$ is assigned based on the image content. Additional details are presented in Algorithm 2. The weight matrix is given prior to the implementation of the algorithm, and there are different generation methods for it across various practical application scenarios. To illustrate, face recognition detection algorithms are used to identify facial regions in datasets such as LFW [37] , and these regions are assigned higher weights. Upon analyzing the following proof, it is evident that, given the same $\epsilon$, the presence of the weight matrix reduces the probability of retaining the original data compared to Algorithm 1, thereby providing stronger privacy protection.
@@ 4.2 Weighted TLDP
    $=\frac{(1-w_{mn})\cdot e^{\epsilon-\frac{I\Delta}{b}}}{2b+w_{mn}\cdot e^{\epsilon-\frac{I\Delta}{b}}}\cdot 2b\cdot e^{\frac{I\Delta}{b}}$
    $\leq\frac{(1-w_{mn})\cdot e^{\epsilon-\frac{I\Delta}{b}}}{2b}\cdot 2b\cdot e^{\frac{I\Delta}{b}}$
    $=(1-w_{mn})\cdot e^{\epsilon}<e^{\epsilon}.$
@@
    $N(x|\sigma^{2})=\frac{1}{\sigma\sqrt{2\pi}}\cdot e^{-\frac{x^{2}}{2\sigma^{2}}},$ (11)
@@
    $\leq\frac{p^{t}[(1-p)\cdot\frac{1}{\sigma\sqrt{2\pi}}]^{I-t}}{[(1-p)\cdot\frac{1}{\sigma\sqrt{2\pi}}\cdot e^{-\frac{\Delta^{2}}{2\sigma^{2}}}]^{I}}$
    $=(\frac{p}{1-p}\cdot\sigma\sqrt{2\pi})^{t}\cdot e^{\frac{I\Delta^{2}}{2\sigma^{2}}}$
@@
    $=\frac{(1-w_{mn})\cdot e^{\epsilon-\frac{I\Delta^{2}}{2\sigma^{2}}}}{\sigma\sqrt{2\pi}+w_{mn}\cdot e^{\epsilon-\frac{I\Delta^{2}}{2\sigma^{2}}}}\cdot\sigma\sqrt{2\pi}\cdot e^{\frac{I\Delta^{2}}{2\sigma^{2}}}$
    $\leq\frac{(1-w_{mn})\cdot e^{\epsilon-\frac{I\Delta^{2}}{2\sigma^{2}}}}{\sigma\sqrt{2\pi}}\cdot\sigma\sqrt{2\pi}\cdot e^{\frac{I\Delta^{2}}{2\sigma^{2}}}$
    $=(1-w_{mn})\cdot e^{\epsilon}<e^{\epsilon}.$
From equation (10) and (12), although the noise reduction achieved by Algorithm 2 is difficult to quantify, the fact that it offers stronger privacy protection than Algorithm 1 is evident as the existence of the term $(1-w_{mn})$.
5 Evaluations
@@
5.1 Setup
5.1.1 Datasets and Tasks
@@
For the computer vision task, we select MNIST [38], CIFAR-10 [39], and SVHN [40] as the datasets. MobileNet (pre-trained on ImageNet)  [^1] https://pytorch.org/vision/main/models/generated/torchvision.models.mobilenet_v2.html#torchvision.models.MobileNet_V2_Weights is used for the computer vision task. The model is trained with the Adam optimizer with a learning rate of 0.001. It uses the cross-entropy loss and has a batch size of 64. We train it for 10 epochs. The MNIST dataset consists of 60,000 training samples and 10,000 test samples of $\times28$ pixel grayscale images of handwritten digits. CIFAR-10 contains 50,000 training samples and 10,000 test samples and has 10 classes of common objects, e.g., birds, airplanes, and cats. SVHN includes door number images collected from Google Street View, and it has 73,257 training images and 26,032 testing images.
@@
5.1.2 Baselines and Metrics
We compare TLDP with other differential privacy mechanisms designed for high-dimensional data. The baselines contain: the Laplace mechanism (Laplace), Gaussian mechanism (Gaussian), Matrix-Variate Gaussian (MVG) [24], Independent Directional Noise (IDN) [25], Layer‐Level Adaptive Gradient Perturbation (LLAGP) [27], and Differential Privacy Adaptive Gradient Clipping Method (DPHSGD) [28]. We implement TLDP-L, TLDP-G, TLDP-L-w, and TLDP-G-w as solutions for applying local differential privacy on private data (see Section 4), where TLDP-L and TLDP-G denote TLDP that uses the Laplace and Gaussian noise, while TLDP-L-w and TLDP-G-w represent TLDP when using Laplace or Gaussian noise while applying a weight matrix for improving utility. Unlike previous studies where noise designed for scalar-valued queries is directly applied to tensors, we mathematically prove that the Laplace and Gaussian noise we add satisfies LDP for tensors (Definition 2), as detailed in Theorem 2. IDN, on the other hand, is a $(\epsilon,\delta)$-differential privacy mechanism for tensor-valued queries. In all experiments, the utility subspace $W$ is set to the identity matrix $E$. LLAGP is designed for adjusting the privacy budget at the layer‐level during training progress, while DPHSGD is for adjusting the clipping threshold of training gradients. Tailored to large-scale data, both of the last two methods work perfectly for conducting comparative tests.
@@ 5.1.2 Baselines and Metrics
$F1_{i}=2\times\frac{P_{i}\times R_{i}}{P_{i}+R_{i}},\\$ (15)
$\textit{F1-score}=\frac{\sum_{i=1}^{n}(w_{i}\times F1_{i})}{\sum_{i=1}^{n}w_{i}},\\$ (16)
5.1.3 Privacy-Preserving Targets
In our experiments, we simulate the distributed computation system using three users and one central server. The users compute their local models using their private dataset and transmit the trained model parameters to the central server. The central server aggregates the received model parameters, computes an average of them, and sends it back to all the users. We denote the models trained with each user’s dataset as local models and the aggregated model on the central server as the global model. Here, we consider three typical tasks in the distributed computation system, i.e., original data, training features, and model parameters. For the first type, users simply transmit their private data to the server, leveraging the enormous computing resources of the server to process and analyze the data. The second type focuses on features extracted from private datasets. This scenario frequently arises in distributed model training across multiple parties, where users process raw data to extract features and transmit the extracted feature embeddings instead of raw data to the central server for model training to avoid privacy leakage. The query mechanism outputs intermediate data/features with differential privacy guarantees, which are then utilized for subsequent training processes on the server side. The third type emphasizes the protection of model parameters. Sharing model parameters is very common in distributed systems, especially when federated learning [43] has evolved for model computation without the need for direct data or feature sharing between data owners and a parameter server. However, research [44] has found that attackers could easily infer the original data used to train the model from intercepted model parameters. Given that the trained model parameters are sensitive, the query function returns a differentially private version of them before transmitting. We present the implementation details categorized by type as follows.
5.2 Implementation Details
5.2.1 Type I: Private Testing Data
The datasets used in Type I include MNIST, CIFAR-10, and SVHN. The setup is illustrated in Table 2. In this type, the training data is likely to be private and should be preserved. Users manipulate their respective data and send it directly to a central server, which uses this information to train the model parameters before transmitting the updated parameters back. No preprocessing or size-trimming is applied to the raw data, ensuring that the accuracy of the experimental results is maintained at the highest possible level.
Additionally, for the MNIST dataset, we include comparisons with the TLDP-L-w and TLDP-G-w methods. In these approaches, the weights are determined based on the grayscale values of the images: higher grayscale values correspond to higher weights. This setting is reasonable since, as the dataset has been preprocessed, apart from the digit regions with values greater than 0, all other parts are set to 0.
@@ 5.2.1 Type I: Private Testing Data
since the range of data is set to (1,1)(-1,1).
Table 2: Datasets and parameter settings for Type I
Dataset Model Training Data Testing Data Batch size $\Delta$
MNIST MobileNet 60000 10000 64 256
CIFAR-10 MobileNet 50000 10000 64 256
SVHN MobileNet 73257 26032 64 256
5.2.2 Type II: Private Training Features
For Type II, we select MNIST, CIFAR-10, and SVHN as experimental datasets. The setup is shown in Table 3. DarkNet and IMDB are not used in this task due to the absence of pre-trained models, and without them, the extracted feature embeddings are prone to being less meaningful. We consider the features extracted from the original training data of these datasets to be confidential and are committed to protecting privacy. To enhance model performance, we substitute the activation function in each layer of MobileNet from ReLU6 to tanh()($\cdot$), ensuring that the output features are normalized within the range of (1,1)(-1,1). MobileNet is divided into two parts: the feature extractor (i.e., all layers except the final fully connected layer) and the classifier (i.e., the final fully connected layer). After extracting features from the feature extractor, noise is added to the features, and only the classifier is trained and tested.